# ALPS II collaboration

### v1

@@ I Introduction
In spite of overwhelming evidence for the necessity of physics beyond the precisely tested Standard Model of particle physics and the technical successes of projects probing the high energy frontier directly, we lack any experimental evidence for an additional energy scale between the electroweak and the Planck scale (see summaries in [ParticleDataGroup:2024cfk]). Hence, interest in accessing higher energy scales indirectly, for example via the search for pseudo-Goldstone bosons, is rising quickly. The most famous example of such a boson is the axion [Peccei:1977hh, Weinberg:1977ma, Wilczek:1977pj], primarily motivated by an explanation of CP conservation in QCD. From astrophysical observations and particle physics experiments, it was quickly noticed that the breaking of the global symmetry giving rise to the axion must occur at very high energies $\mathrm{}{f_{a}}$ above the electroweak scale (see [Sikivie:1983ip] and references therein). This axion was named “invisible” as its coupling strengths to Standard Model constituents are predicted to be proportional to $\mathrm{}{f_{a}^{-1}}$. While $\mathrm{}{f_{a}}$ being much larger than the electroweak scale comes with huge experimental challenges, it makes the axion an ideal cold dark matter candidate [Abbott:1982af, Preskill:1982cy, Dine:1982ah, Arias:2012az, DiLuzio:2020wdo, Chadha-Day:2021szb, Eroncel:2022vjg], adding strong cosmological motivation to the particle physics science case. Furthermore, axion-like particles have been proposed to explain a small cosmological constant and generate a naturally small electroweak scale [Graham:2019bfu, Espinosa:2015eda]. Recent results from the DESI large-scale cosmological survey have again sparked discussions on axion dark energy (see for example [Seto:2024cgo, Qu:2024lpx]).
In general, many theories for physics beyond the Standard Model predict the existence of axions, axion-like particles, hidden photons, or other WISPs (Weakly Interacting Slim Particles, introduced by A. Ringwald in 2007 at the 3rd Joint ILIAS–CERN–DESY Axion–WIMPs workshop) [Arvanitaki:2009fg, Cicoli:2012sz, Jaeckel:2010ni, Hui:2021tkt, Dvali:2024zpc]. Due to their extremely weak interactions, such WISPs cannot be found at existing or planned accelerator-based experiments. Instead, dedicated setups are needed, which are reviewed in [Sikivie:2020zpn]. World-wide, three approaches are followed:
@@
The Any-Light-Particle-Search II (ALPS II) experiment at DESY in Hamburg is a prime example for the third category, one which is independent of astrophysical and cosmological assumptions. Within its particle physics strategy, DESY also plans to host the helioscope BabyIAXO [IAXO:2020wwp] and the haloscope MADMAX [MADMAX:2024jnp, MADMAX:2024sxs].
ALPS II follows the light-shining-through-a-wall (LSW) approach [Redondo:2010dp, Arias:2010bh]:
@@
ALPS II builds upon experiences with ALPS [ALPS:2009des, Ehret:2010mh], the first LSW experiment at DESY. At design sensitivity, it will surpass the axion-photon coupling-strength sensitivity of OSQAR at CERN [OSQAR:2015qdv], the previously most sensitive LSW-experiment, by three orders of magnitude.
@@ II Boson detection in LSW experiments
LSW experiments provide probabilities (or upper limits) for photons to seemingly pass light-tight walls. In the following, we briefly outline how to determine properties of scalar and pseudoscalar, vector and tensor bosons from such measurements. Using ALPS II data to look for minicharged particles like in [Ehret:2010mh] or high-frequency gravitational waves (compare [Ejlli:2019bqj, Ringwald:2020ist]) as well as more complex scenarios involving several axion-like particles will be addressed in future publications.
@@ II.1 Scalar and pseudoscalar bosons
We follow references [Redondo:2010dp, Arias:2010bh], where the interaction of scalar and pseudoscalar bosons with electromagnetic fields is described by effective Lagrangians including
@@ II.2 Vector bosons
Massive vector bosons $\mathrm{}{\gamma^{\prime}}$ (frequently named “hidden” or “dark photons”) do not require any magnetic field for mixing with photons. Following again [Arias:2010bh], one derives, with the kinetic mixing parameter $\mathrm{}{\epsilon}$ and the hidden vector potential $\mathrm{}{X_{\mu}}$, from
@@ II.3 Tensor bosons
The interaction of photons with massive parity-even spin-2 bosons $\chi$ in a magnetic field is worked out in [Garcia-Cely:2025ula]:
@@ II.4 Standard model background
An LSW effect can be mediated by photon-graviton conversion in a background magnetic field. With Newton’s constant $G$, this probability is estimated as (see [Raffelt:1987im] and [Krizova:2024lvy], eq. 3.157):
@@
At the ALPS II design sensitivity [Bahre:2013ywa], this results in a power of $\mathrm{}{\approx 4\cdot 10^{-54}}$ W behind the wall, which is totally negligible for any practical purposes.
@@ III The ALPS II experiment
The experiment ALPS II is located in a straight section of the former HERA accelerator complex [19436] around the hall North at DESY in Hamburg. ALPS II, as laid out in [Bahre:2013ywa], is striving for the first realization of an optical resonantly enhanced LSW experiment proposed earlier [Hoogeveen:1990vq, Fukuda:1996kwa, Sikivie:2007qm]. It consists of two strings of modified superconducting dipole magnets (originally built for the HERA proton beam) before and after the wall as well as three cleanrooms at both ends and in the center to host the optical components.
@@ III.1 The magnet strings
The magnets and the infrastructure of HERA, shut down in 2007, provide a unique environment for a large LSW experimental setup, which was first noticed in [Ringwald:2003nsa]. ALPS II relies especially on the following opportunities:
@@
  • Infrastructure and the cryogenics for cooling the magnets [Horlitz:1984ym] could be re-used with limited modifications.
@@
However, the cold mass inside an unmodified HERA dipole magnet is bent as the magnet was constructed to guide protons through the arcs of the accelerator. This would have limited the horizontal free aperture to 35 mm, insufficient for a high-finesse optical cavity making full use of a HERA straight section. Therefore, a “brute-force” straightening procedure was developed [Albrecht:2020ntd]. Without removing the cold mass from the cryostat, suspensions of the cold mass were changed and special pressure-props installed at the positions of cryostat flanges to recover, on average, a 50 mm horizontal aperture. It is remarkable that the straightened dipoles show essentially the same performance specification in tests as at the time of their fabrication about 40 years ago [Albrecht:2020ntd]. This success allowed us to install 12 plus 12 straightened magnets in the straight tunnel section around the HERA North hall. The parameters of the magnet strings are summarized in Table 1. Note that the magnetic lengths in front of and behind the wall are equal: $\mathrm{}{L=N\cdot L_{\mathrm{SD}}}$.
@@ III.2 The optical system
While the final optical system of ALPS II will feature two mode-matched optical resonators [Ortiz:2020tgs], we only implemented the regeneration cavity (RC) behind the wall [Kozlowski:2024jzm] for the first science run. This was done to simplify the operation and to increase potential stray-light “leak” intensities through the wall by a factor of 40 so that they could be traced more easily. Fig. 1 sketches the optics system. Details of the system, its calibration and data analyses will be presented in a forthcoming publication [aaron].
Figure 1: The optical system implemented in the first science run of ALPS II (taken from [aaron]). “FSR” denotes the free spectral range of the regeneration cavity, “North Left”, “Central Hall” and “North Right” refer to the three cleanrooms at the ALPS II magnet strings. The black bar on the “central optical bench” indicates the light-tight wall housing a shutter which is opened for system checks and calibration. See the text for a brief explanation and [aaron] for details. Twelve HERA dipole magnets each surround the production area and the regeneration cavity.
@@
The frequency of the local oscillator laser (blue in Fig. 1) is again stabilized with respect to the auxiliary laser with a frequency offset such that it is not resonant with the regeneration cavity. The local oscillator laser light reflected from the cavity interferes with the light (potentially) regenerated after the wall that is leaving the cavity as it transmits the mirror RC1. A heterodyne sensing scheme [aaron, Hallal:2020ibe] is used to sense the regenerated power via a measurement of the interference beatnote between these two fields at the photodetector $\mathrm{PD_{science}}$. If a signal were to appear, the beatnote frequency would be the known fixed frequency difference between the local oscillator laser and the high-power laser, referred to here as the heterodyne frequency $f_{\mathrm{s}}$.
The most relevant parameters of the optical system used here are given in Table 2. We note that the collaboration has developed an alternative optical scheme using a Transition Edge Sensor detector system [Gimeno:2023nfr] to count single photons generated behind the wall. This system could be implemented at ALPS II to cross-check a discovery with the heterodyne sensing scheme described here.
@@
Table 2: Parameters of the optical system. $\mathrm{}{P_{\mathrm{i}}}$ denotes the laser power traversing the magnet string before the wall. $\mathrm{}{P_{\mathrm{i}}}$, $\beta$ and $|\eta|^{2}$ varied during data taking (see [aaron]). $L_{\mathrm{RC}}$ is measured via the free spectral range of the cavity.
@@ IV Data analysis
A shutter in the light-tight wall is used to help calibrate ALPS II. It can be opened to allow a tiny fraction of the HPL light to enter the regeneration cavity. Then, the probability of a photon-WISP conversion $\mathrm{}{\mathcal{P}_{\gamma\leftrightarrow\mathrm{WISP}}}$ is experimentally determined by comparing the powers at the HPL frequency reaching the science photodetector $\mathrm{{PD_{science}}}$ (behind the cavity mirror $\mathrm{{{RC1}}}$ in Fig. 1) with the shutter in the wall open and closed. The powers are determined by evaluating the heterodyne beat signal. The procedure is sketched below. Details can be found in [aaron]. We use the following notations for light powers at the HPL frequency:
@@ IV.2 Closed-shutter mode
The re-conversion of bosons to light is resonantly enhanced by the cavity [Hoogeveen:1990vq, Fukuda:1996kwa, Sikivie:2007qm]. This light generated inside the cavity leaves the cavity through one of its mirrors, or otherwise through a cavity loss mechanism. The resonant enhancement factor for light generated inside the cavity and sensed behind $\mathrm{{{RC1}}}$ is approximately
@@
Most systematic uncertainties are identical in $\mathrm{}{P_{\gamma}}$ and $\mathrm{}{P_{\mathrm{open}}}$ and hence cancel in the ratio. Again, details on corresponding cross-checks are given in [aaron].
@@ IV.3 Experimental results
The main experimental results are exemplified in Fig. 2, which displays the spectral power normalized by the open shutter power, $\mathrm{}{P_{\gamma}(\Delta f_{\mathrm{s}})/P_{\mathrm{open}}}$, for frequency offsets $\Delta f_{\mathrm{s}}$ from the signals’s heterodyne frequency $f_{\mathrm{s}}$. These normalized spectral power distributions were determined by evaluating the power modulation on $\mathrm{{PD_{science}}}$ at frequencies in the vicinity of $f_{\mathrm{s}}$. A clear signal above background shows up at $\Delta f_{\mathrm{s}}=0\,\mathrm{Hz}$. However, this peak is much broader than expected for LSW mediated by new bosons. It is likely caused by stray-light passing around the wall (see [aaron]) with an intensity of a few $\mathrm{}{10^{-22}}$ W at the HPL frequency.
Figure 2: Taken from [aaron]: normalized spectral power from measurements on $\mathrm{{PD_{science}}}$ determined for frequencies around the signal’s heterodyne frequency for the polarization $\mathrm{}{\gamma_{\parallel}}$ (Table 3). An LSW signal would only show up in the frequency bin centered at 0 Hz. The observed broader distribution is likely caused by stray-light.
In Ref. [aaron], the background at the signal frequency is estimated using the measured background at nearby frequencies. The ratios of the results at the signal frequency to the estimated backgrounds are shown in column 3 of Table 3. The results agree well with the expectation value of 1 in the absence of a signal. In column 4 of the same table, we derive 95% confidence level (CL) upper limits for the ratio $P_{\gamma}\,/\,P_{\mathrm{open}}$. In Ref. [aaron], it is shown that, at frequencies other than the signal frequency, the results adhere to a non-central $\mathrm{\chi^{2}}$ distribution. The limits are therefore based on the assumption that the results at the signal frequency show a similar behavior. Here, we define a 95% CL as the probability to have not missed a detection corresponding to a significance of five standard deviations. Using the measurement of $\mathrm{}{T_{\mathrm{COB}}T_{\mathrm{RC2}}}=(9.7\pm 1.2)\cdot 10^{-23}$, limits on the conversion probabilities $\mathrm{{\gamma\rightarrow WISP\rightarrow\gamma}}$ can be determined. Following [COUSINS1992331], we take a conservative approach and add the systematic uncertainties to the limits derived from statistical uncertainties, resulting in the limits on the conversion probabilities in column 4.
@@
Table 3: The main experimental results of the ALPS II first science campaign. “Polarization” is defined with respect to the orientation of the magnetic field. In the third column the power ratio of the closed- and open-shutter measurements at the HPL frequency is compared to a background estimation as explained in [aaron]. Here, the errors are statistical uncertainties as most of the systematic uncertainties cancel. The fourth column displays upper limits on the power ratio of the closed- and open-shutter measurements with their systematic uncertainties. The last column gives the upper limits on boson-photon conversion probabilities incorporating the systematic uncertainties (p. 2). Note also our definition of the 95% CL on page 2.

### v2

@@ I Introduction
In spite of overwhelming evidence for the necessity of physics beyond the precisely tested Standard Model of particle physics and the technical successes of projects probing the high energy frontier directly, we lack any experimental evidence for an additional energy scale between the electroweak and the Planck scale (see summaries in [44]). Hence, interest in accessing higher energy scales indirectly, for example via the search for pseudo-Goldstone bosons, is rising quickly. The most famous example of such a boson is the axion [47, 63, 64], primarily motivated by an explanation of CP conservation in QCD. From astrophysical observations and particle physics experiments, it was quickly noticed that the breaking of the global symmetry giving rise to the axion must occur at very high energies $\mathrm{}{f_{a}}$ above the electroweak scale (see [59] and references therein). This axion was named “invisible” as its coupling strengths to Standard Model constituents are predicted to be proportional to $\mathrm{}{f_{a}^{-1}}$. While $\mathrm{}{f_{a}}$ being much larger than the electroweak scale comes with huge experimental challenges, it makes the axion an ideal cold dark matter candidate [1, 48, 21, 6, 19, 15, 28], adding strong cosmological motivation to the particle physics science case. Furthermore, axion-like particles have been proposed to explain a small cosmological constant and generate a naturally small electroweak scale [33, 29]. Recent results from the DESI large-scale cosmological survey have again sparked discussions on axion dark energy (see for example [56, 49]).
In general, many theories for physics beyond the Standard Model predict the existence of axions, axion-like particles, hidden photons, or other WISPs (Weakly Interacting Slim Particles, introduced by A. Ringwald in 2007 at the 3rd Joint ILIAS–CERN–DESY Axion–WIMPs workshop) [8, 16, 39, 38, 22]. Due to their extremely weak interactions, such WISPs cannot be found at existing or planned accelerator-based experiments. Instead, dedicated setups are needed, which are reviewed in [60]. World-wide, three approaches are followed:
@@
The Any-Light-Particle-Search II (ALPS II) experiment at DESY in Hamburg is a prime example for the third category, one which is independent of astrophysical and cosmological assumptions. Within its particle physics strategy, DESY also plans to host the helioscope BabyIAXO [2] and the haloscope MADMAX [23, 31].
ALPS II follows the light-shining-through-a-wall (LSW) approach [51, 7]:
@@
ALPS II builds upon experiences with ALPS [24, 25], the first LSW experiment at DESY. At design sensitivity, it will surpass the axion-photon coupling-strength sensitivity of OSQAR at CERN [11], the previously most sensitive LSW-experiment, by three orders of magnitude.
@@ II Boson detection in LSW experiments
LSW experiments provide probabilities (or upper limits) for photons to seemingly pass light-tight walls. In the following, we briefly outline how to determine properties of scalar and pseudoscalar, vector and tensor bosons from such measurements. Using ALPS II data to look for minicharged particles like in [25] or high-frequency gravitational waves (compare [27, 54]) as well as more complex scenarios involving several axion-like particles [18] will be addressed in future publications.
@@ II.1 Scalar and pseudoscalar bosons
We follow references [51, 7], where the interaction of scalar and pseudoscalar bosons with electromagnetic fields is described by effective Lagrangians including
@@ II.2 Vector bosons
Massive vector bosons $\mathrm{}{\gamma^{\prime}}$ (frequently named “hidden” or “dark photons”) do not require any magnetic field for mixing with photons. Following again [7], one derives, with the kinetic mixing parameter $\mathrm{}{\epsilon}$ and the hidden vector potential $\mathrm{}{X_{\mu}}$, from
@@ II.3 Tensor bosons
For the interaction of photons with massive parity-even spin-2 bosons $\chi$ in a magnetic field, we follow the formalism of [32]:
@@ II.4 Standard model background
An LSW effect can be mediated by photon-graviton conversion in a background magnetic field. With Newton’s constant $G$, this probability is estimated as (see [50] and [43], eq. 3.157):
@@
At the ALPS II design sensitivity [10], this results in a power of $\mathrm{}{\approx 4\cdot 10^{-54}}$ W behind the wall, which is totally negligible for any practical purposes.
@@ III The ALPS II experiment
The experiment ALPS II is located in a straight section of the former HERA accelerator complex [35] around the hall North at DESY in Hamburg. ALPS II, as laid out in [10], is striving for the first realization of an optical resonantly enhanced LSW experiment proposed earlier [36, 30, 58]. It consists of two strings of modified superconducting dipole magnets (originally built for the HERA proton beam) before and after the wall as well as three cleanrooms at both ends and in the center to host the optical components.
@@ III.1 The magnet strings
The magnets and the infrastructure of HERA, shut down in 2007, provide a unique environment for a large LSW experimental setup, which was first noticed in [53]. ALPS II relies especially on the following opportunities:
@@
  • Infrastructure and the cryogenics for cooling the magnets [37] could be re-used with limited modifications.
@@
However, the cold mass inside an unmodified HERA dipole magnet is bent as the magnet was constructed to guide protons through the arcs of the accelerator. This would have limited the horizontal free aperture to 35 mm, insufficient for a high-finesse optical cavity making full use of a HERA straight section. Therefore, a “brute-force” straightening procedure was developed [4]. Without removing the cold mass from the cryostat, suspensions of the cold mass were changed and special pressure-props installed at the positions of cryostat flanges to recover, on average, a 50 mm horizontal aperture. It is remarkable that the straightened dipoles show essentially the same performance specification in tests as at the time of their fabrication about 40 years ago [4]. This success allowed us to install 12 plus 12 straightened magnets in the straight tunnel section around the HERA North hall. The parameters of the magnet strings are summarized in Table 1. Note that the magnetic lengths in front of and behind the wall are equal: $\mathrm{}{L=N\cdot L_{\mathrm{SD}}}$.
@@ III.2 The optical system
While the final optical system of ALPS II will feature two mode-matched optical resonators [46], we only implemented the regeneration cavity (RC) behind the wall [42] for the first science run. This was done to simplify the operation and to increase potential stray-light “leak” intensities through the wall by a factor of 40 so that they could be traced more easily. Fig. 1 sketches the optics system. Details of the system, its calibration and data analyses will be presented in a forthcoming publication [62].
Figure 1: The optical system implemented in the first science run of ALPS II (taken from [62]). “FSR” denotes the free spectral range of the regeneration cavity, “North Left”, “Central Hall” and “North Right” refer to the three cleanrooms at the ALPS II magnet strings. The black bar on the “central optical bench” indicates the light-tight wall housing a shutter which is opened for system checks and calibration. See the text for a brief explanation and [62] for details. Twelve HERA dipole magnets each surround the production area and the regeneration cavity.
@@
The frequency of the local oscillator laser (blue in Fig. 1) is again stabilized with respect to the auxiliary laser with a frequency offset such that it is not resonant with the regeneration cavity. The local oscillator laser light reflected from the cavity interferes with the light (potentially) regenerated after the wall that is leaving the cavity as it transmits the mirror RC1. A heterodyne sensing scheme [62, 34] is used to sense the regenerated power via a measurement of the interference beatnote between these two fields at the photodetector $\mathrm{PD_{science}}$. If a signal were to appear, the beatnote frequency would be the known fixed frequency difference between the local oscillator laser and the high-power laser, referred to here as the heterodyne frequency $f_{\mathrm{s}}$.
The most relevant parameters of the optical system used here are given in Table 2. We note that the collaboration has developed an alternative optical scheme using a Transition Edge Sensor detector system [55] to count single photons generated behind the wall. This system could be implemented at ALPS II to cross-check a discovery with the heterodyne sensing scheme described here.
@@
Table 2: Parameters of the optical system. $\mathrm{}{P_{\mathrm{i}}}$ denotes the laser power traversing the magnet string before the wall. $\mathrm{}{P_{\mathrm{i}}}$, $\beta$ and $|\eta|^{2}$ varied during data taking (see [62]). $L_{\mathrm{RC}}$ is measured via the free spectral range of the cavity.
@@ IV Data analysis
A shutter in the light-tight wall is used to help calibrate ALPS II. It can be opened to allow a tiny fraction of the HPL light to enter the regeneration cavity. Then, the probability of a photon-WISP conversion $\mathrm{}{\mathcal{P}_{\gamma\leftrightarrow\mathrm{WISP}}}$ is experimentally determined by comparing the powers at the HPL frequency reaching the science photodetector $\mathrm{{PD_{science}}}$ (behind the cavity mirror $\mathrm{{{RC1}}}$ in Fig. 1) with the shutter in the wall open and closed. The powers are determined by evaluating the heterodyne beat signal. The procedure is sketched below. Details can be found in [62]. We use the following notations for light powers at the HPL frequency:
@@ IV.2 Closed-shutter mode
The re-conversion of bosons to light is resonantly enhanced by the cavity [36, 30, 58]. This light generated inside the cavity leaves the cavity through one of its mirrors, or otherwise through a cavity loss mechanism. The resonant enhancement factor for light generated inside the cavity and sensed behind $\mathrm{{{RC1}}}$ is approximately
@@
Most systematic uncertainties are identical in $\mathrm{}{P_{\gamma}}$ and $\mathrm{}{P_{\mathrm{open}}}$ and hence cancel in the ratio. Again, details on corresponding cross-checks are given in [62].
@@ IV.3 Experimental results
The main experimental results are exemplified in Fig. 2, which displays the spectral power normalized by the open shutter power, $\mathrm{}{P_{\gamma}(\Delta f_{\mathrm{s}})/P_{\mathrm{open}}}$, for frequency offsets $\Delta f_{\mathrm{s}}$ from the signals’s heterodyne frequency $f_{\mathrm{s}}$. These normalized spectral power distributions were determined by evaluating the power modulation on $\mathrm{{PD_{science}}}$ at frequencies in the vicinity of $f_{\mathrm{s}}$. A clear signal above background shows up at $\Delta f_{\mathrm{s}}=0\,\mathrm{Hz}$. However, this peak is much broader than expected for LSW mediated by new bosons. It is likely caused by stray-light passing around the wall (see [62]) with an intensity of a few $\mathrm{}{10^{-22}}$ W at the HPL frequency.
Figure 2: Taken from [62]: normalized spectral power from measurements on $\mathrm{{PD_{science}}}$ determined for frequencies around the signal’s heterodyne frequency for the polarization $\mathrm{}{\gamma_{\parallel}}$ (Table 3). An LSW signal would only show up in the frequency bin centered at 0 Hz. The observed broader distribution is likely caused by stray-light.
In Ref. [62], the background at the signal frequency is estimated using the measured background at nearby frequencies. The ratios of the results at the signal frequency to the estimated backgrounds are shown in column 3 of Table 3. The results agree well with the expectation value of 1 in the absence of a signal. In column 4 of the same table, we derive 95% confidence level (CL) upper limits for the ratio $P_{\gamma}\,/\,P_{\mathrm{open}}$. In Ref. [62], it is shown that, at frequencies other than the signal frequency, the results adhere to a non-central $\mathrm{\chi^{2}}$ distribution. The limits are therefore based on the assumption that the results at the signal frequency show a similar behavior. Here, we define a 95% CL as the probability to have not missed a detection corresponding to a significance of five standard deviations. Using the measurement of $\mathrm{}{T_{\mathrm{COB}}T_{\mathrm{RC2}}}=(9.7\pm 1.2)\cdot 10^{-23}$, limits on the conversion probabilities $\mathrm{{\gamma\rightarrow WISP\rightarrow\gamma}}$ can be determined. Following [17], we take a conservative approach and add the systematic uncertainties to the limits derived from statistical uncertainties, resulting in the limits on the conversion probabilities in column 4.
@@
Table 3: The main experimental results of the ALPS II first science campaign. “Polarization” is defined with respect to the orientation of the magnetic field. In the third column the power ratio of the closed- and open-shutter measurements at the HPL frequency is compared to a background estimation as explained in [62]. Here, the errors are statistical uncertainties as most of the systematic uncertainties cancel. The fourth column displays upper limits on the power ratio of the closed- and open-shutter measurements with their systematic uncertainties. The last column gives the upper limits on boson-photon conversion probabilities incorporating the systematic uncertainties (p. 2). Note also our definition of the 95% CL on page 2.